OpenStudy (babalooo):

confused... if I had the equation x''+x=0 subjected to the conditions of x=pi/12 and x'=0, would I get x''=0 or x''=-pi/12? derivative of x'=x''=0 but x''=-x which means x''=-pi/12... so which is correct...

OpenStudy (babalooo):

@ganeshie8

OpenStudy (anonymous):

your last statement is correct....

OpenStudy (anonymous):

for the x''+x=0 we have the general solution given by x = Acost + B sint where A,B are the constants which we need to find

OpenStudy (anonymous):

u may check what are the initail conditions...

OpenStudy (anonymous):

Equation 1 might be... \[A\cos t+B\sin t=\dfrac{\pi}{12}\] Equation 2... \[-A\cos t+B\sin t=0\] Solve for A and B...

OpenStudy (anonymous):

oops wrong Equation 2...

OpenStudy (anonymous):

it should be... \[-A \sin t+B\cos t=0\]

OpenStudy (anonymous):

@matricked thanks for your input...